The table below provides information for the economy of Zawi.
 

C = 25 + 0.6YXN = 15 − 0.1Y
I = 100G = 100


a. The value of equilibrium income is $ 

b. Set up a balancing row to verify your calculations (the tax equation is T = 80 + 0.15Y and X = 210). Enter your responses as whole numbers.
 

YTYDCSIGXIMXNAE

 
c. If exports decrease by 60, the new equilibrium income is $ 
Answer #1

ANSWER :


a.


Equilibrium income, Y 

= C + I + G + X net 

= 25 + 0.6Y + 100 + 100 + 15 - 0.1y

= 240 + 0.5Y

=> Y - 0.5Y = 240

=> 0.5Y = 240 

=> Y = 480 (ANSWER)


b. 


T = 80 + 0.15Y = 80 + 0.15*480 = 152 

C = 25 + 0.6Y = 25 + 0.6*480 = 313

X net = 15 - 0.1Y = 15 - 0.1*480 = - 33 

=> IM = X - X net = 210 - (-33) = 243 



  Y        T         Yd          C          S           I           G             X         IM


480    152      328       313        15       100        100          210       243   



Check : Y = C + I + G + X - IM = 313 + 100 + 100 + 210 - 243 = 480 (OK)


c. 


If X decreases by 60, Y also decrease by 60.

So, new equilibrium income = 480 - 60 = 420 (ANSWER).

answered by: Tulsiram Garg